OpenStudy (anonymous):

How do you solve tan(v) if sin(v)=1/3?

OpenStudy (anonymous):

we just think of the domain of v is from 0 to 2pi use \[v=\sin ^{-1}(1/3)\] we can have the value of v then work tan v u can do it on your calculator

OpenStudy (anonymous):

Have to do this without the help of calculator.. :p

OpenStudy (anonymous):

tanv=sinv/cosv

OpenStudy (goformit100):

use basic trigonometrical formulae

OpenStudy (anonymous):

tan(v)*cos(v)=1/3.. then what?

OpenStudy (anonymous):

ok, (sinv)^2+(cosv)^2=1 sinv=1/3 so (cosv)^2=1-(1/3)^2=8/9 \[cosv =2\sqrt{2}/3\] and tan v=(sinv)/(cosv) it should have TWO answers!

OpenStudy (anonymous):

got it?

OpenStudy (anonymous):

\[cosv=\pm 2\sqrt{2}/3\]

OpenStudy (phi):

to answer these questions I draw a triangle, and label the known sides. use pythagoras to fill in the unknown side.